How do you find a triangle whose ratio is 2:5:8 - this is sophmore advanced geometry and I am just a homebound teacher. please help!!!
Answer by checkley77(12844)

2:5:8
IF these are the ratios of the three sides.
2n+5n+8n=15n where n is the multiplier of the sides..
However; these ratios are not valid for a triangle.
The sum of the 2 smallest sides (2+5)=7 is less than the third side=8.
Thus these ratios cannot be used to make a triangle.
-----------------------------------------------------------
IF these ratios are for the angles then:
2n+5n+8n=15n
15n=180
n=180/15
n=12 is the multiplier of the ratio values.
2*12=24 degrees.
5*12=60 degrees.
8*12=96 degrees.
Proof:
24+60+96=180
180=180